MATLAB question: I have some data on excell and I have to write a code that does the following:

Read from Data_all.xlsx Sheet “Wave_data, your specific Wave Height (WH) measurements and using a for loop and if-elseif-
Please use dummy data for your convenience.

Thank you in advance for your help!
Answer #1

Here is the MATLAB code to perform the specified task.

 clear; clc; FileName = 'Data_all' %Keep the file in the same folder as .m file num = xlsread(FileName); %Read data from excel sheet %Initialize counters a=0; b=0; c=0; d=0; e=0; f=0; g=0; for i = 1:1:length(num) if num(i) == 0 | num(i) == 1 a = a+1; elseif num(i) == 2 | num(i) == 3 b = b+1; elseif num(i) == 4 | num(i) == 5 c = c+1; elseif num(i) == 6 | num(i) == 7 d = d+1; elseif num(i) == 8 | num(i) == 9 e=e+1; elseif num(i) == 10 | num(i) == 11 f = f+1; elseif num(i) == 12 | num(i) == 13 g =g+1; else disp("Invalid Input"); end end y = [ a b c d e f g]; name = {'0-1';'2-3';'4-5';'6-7';'8-9'; '10-11'; '12-13'}; x = [1:7] %For X- axis %figure(1); %Uncomment these two lines for having subplot in your next task %subplot(2,1,1); bar(x,y) %Plot Bar graph xlabel('Wave height in meters'); %Label on X-axis ylabel('Number of waves'); %Label in Y-axis set(gca,'xticklabel',name); %Set wave height for each bar in graph

The following screenshot will help you with the indentation:

TII IIIIIIIIIIIIIIIIIIIIIIIIIII clear; clc; FileName = Data_all Keep the file in the same folder as .m file num = xlsread (

After giving the random numbers from 0 - 13 from excel sheet as follows:

The output will be as follows:

- OX Figure 1 Eile Edit View Insert Tools Desktop Window Help Number of waves 0-1 2-3 10-11 12-13 4-5 6-7 8-9 Wave height in

Uncomment the figure(1) and subplot(2,1,1); for subplot in your next task.

Note: Feel free to reach me out using the comment section if needed help. Please give a thumbs up for the solution. Thanks in advance.